Comment lampoly SNF 2009-07-12 23:53:33: Gap check

mtang at snf.stanford.edu mtang at snf.stanford.edu
Sun Jul 12 23:53:33 PDT 2009


Checked gap on Friday.  It looks like it was off by about 6 mm.  Not significant for the process, but possibly so for the hardware (thus, the recent damage.)
Gap Cal parameter index 256 was changed from 7116 to 6496 (based on 0.62 cm delta).
Hard stops still need to be adjusted, especially the upper ones.  Mechanical switch position for upper and lower limits need to be adjusted as well.
After this, the new ring clamp with clamp ceramics can be installed (hopefully, the ceramics will arrive Monday.)
Entrance loadlock door was left open to make it easier to see.   Please make sure to manually shut it before pumping down the chamber.
Also the electrode temperature needs to be set back up to 60 C.  After the installation of the ring and system pump/purge cycle, rate of rise should be done.
If all looks good, 20 wafers should be run on recipe 99 and then recipe 1.    The He backside flow should be checked:  at 6 torr, the flow should be stable at 8-10 sccm.  (It may take several wafers to get there.)
If all looks good, then an etch rate test/qual should be done.




More information about the lampoly-pcs mailing list